# Domain Generalizable Knowledge Tracing via Concept Aggregation and Relation-Based Attention

### v1

@@ III-B Model Architecture
[…]
@@
Knowledge tracing backbone. To encode question embedding and question-response embedding into knowledge state, we employ backbone of existing knowledge tracing model. Here we take DKT, SAINT and AKT as examples to illustrate how to transform embedding into knowledge state using off-the-shelf knowledge tracing models.
For DKT, we use LSTM as knowledge state emcoder to get student’s knowledge state:
$\begin{split}f_{t}&=\sigma(W_{f}e_{qr_{t}}+U_{f}h_{t-1}+b_{f}),\\ i_{t}&=\sigma(W_{i}e_{qr_{t}}+U_{i}h_{t-1}+b_{i}),\\ o_{t}&=\sigma(W_{o}e_{qr_{t}}+U_{o}h_{t-1}+b_{o}),\\ \tilde{c_{t}}&=\sigma(W_{c}e_{qr_{t}}+U_{c}h_{t-1}+b_{c}),\\ c_{t}&=f_{t}\odot c_{t-1}+i_{t}\odot\tilde{c_{t}},\\ h_{t}&=o_{t}\cdot\sigma(c_{t}),\end{split}$ (3)
where $e_{qr_{t}}$ denotes the t-th question-response feature and $h_{t}$ represents the student’s knowledge state at timestep t.
Likewise, in the AKT model, three attention-based modules are utilized: the question encoder, the knowledge encoder, and the knowledge retriever. The student’s knowledge state can be represented as:
$\begin{split}x_{t}&=\text{Encoder}_{q}(e_{q_{1}},e_{q_{2}},...,e_{q_{t}}),\\ y_{t}&=\text{Encoder}_{k}(e_{qr_{1}},e_{qr_{2}},...,e_{qr_{t}}),\\ h_{t}&=\text{Decoder}((x_{1},y_{1}),(x_{2},y_{2}),...,(x_{t-1},y_{t-1}),x_{t}),\end{split}$ (4)
where $\text{Encoder}_{q}$ represents the question encoder which produce modified, contextualized representations of each question based on the sequence of questions the learner has previously practiced on, $\text{Encoder}_{k}$ represents the knowledge encoder which produces modified, contextualized representations of the knowledge the learner acquired while responding to past questions, and Decoder represents the knowledge retriever which retrieves knowledge acquired in the past that is relevant to the current question using an attention mechanism. In AKT, the student’s knowledge state $h$ is obtained in the knowledge retriever module by first constructing question features $x$ based on the student’s question sequence and constructing question-interaction features $y$ based on the student’s question-interaction sequence. Then, $h$ is derived through attention layers applied to $x$ and $y$.
As for SAINT, the transformer’s encoder is responsible for receiving the student’s question sequence, while the decoder receives both the student’s question-interaction sequence and the output from the encoder:
$\begin{split}o_{t}&=\text{Encoder}(e_{q_{1}},e_{q_{2}},...,e_{q_{t}}),\\ h_{t}&=\text{Decoder}(e_{qt_{1}},e_{qt_{2}},...,e_{qt_{t-1}},o_{1},o2,...,o_{t}),\end{split}$ (5)
where Encoder and Decoder represent the Transformer’s encoder and decoder respectively.
Knowledge state decoder. To decode knowledge state into the probability, we use the same knowledge encoder to predict the probability of a student answering the next question correctly :
@@
where Dec represents the knowledge state decoder, $W_{1}$, $W_{2}$ and $b_{1}$, $b_{2}$ denote weights and biases, respectively. $\sigma(\cdot)$ is the sigmoid function.
III-C Sequence Instance Normalization
To reduce the distribution discrepancy from different domains, utilizing normalization method is a common solution. However, existing normalization methods may not be applicable for sequence feature in knowledge tracing task since feature after timestep $t$ should be kept unseen when calculating $y$ at timestep $t$. To tackle the dilemma, we design the sequence instance normalization SeqIN in our KT model to normalize the feature embeddings of sequential student interactions among domains. Since the normalization process takes into account the fact that the later interactions cannot be seen in the previous interactions, the feature embedding of the interaction each moment is normalized by only the statistics of all the previous moments.
Given a sequential feature matrix $M=(m_{1},m_{2},...,m_{n})$ where $m_{t}\in\mathbb{R}^{d}$ is the feature embedding at time $t$, the normalized feature matrix $\tilde{M}$ is calculated by
$\tilde{M}=(\tilde{m}_{1},\tilde{m}_{2},...,\tilde{m}_{n}),$ (7)
$\tilde{m}_{t}=\gamma(\frac{m_{t}-\mu_{t}(M)}{\sigma_{t}(M)})+\beta,$ (8)
$\mu_{t}(M)=\frac{1}{t+1}(p+\sum_{i=1}^{t}m_{i}),$ (9)
$\sigma_{t}(M)=\sqrt{\frac{(p-\mu_{t}(M))^{2}+\sum_{i=1}^{t}(m_{i}-\mu_{t}(M))^{2}}{t+1}},$ (10)
where $\gamma,\beta\in\mathbb{R}^{d}$ are the parameters learned from the data. $\mu_{t}(M)$ and $\sigma_{t}(M)$ denote the mean and standard deviation of the sequence $\{m_{1},m_{2},...,m_{t}\}(1\leq t\leq n)$. As seen, SeqIN normalizes $m_{t}$ by considering all previous feature embeddings up to time $t$ while remaining later feature embeddings unseen. In consideration of the meaninglessness to compute the mean and standard deviation of $m_{1}$ itself, we add a padding vector $p$ in front of the original sequence, i.e., $\{p,m_{1},m_{2},...,m_{n}\}$, where $p$ is learned along with the model. Moreover, SeqIN can allow the feature embedding matrices of student interactions from different source domains to be aggregated together, which will be obviously observed in Fig. 5 in the visualization experiments.
We apply SeqIN to aforementioned knowledge tracing backbone networks, including DKT, AKT and SAINT. For DKT, the SeqIN is directly used on students’ knowledge state h:
@@ III-C Sequence Instance Normalization
where $h$ is the students’ knwoledge state. For AKT, we apply SeqIN to intermediate features x and y as follows:
@@
As for SAINT, we apply SeqIN to intermediate features o:
$\tilde{o}=SeqIN(o),$ (14)
where $o$ is the output of Transformer’s encoder.
III-D Concept Aggregation
We here illustrate the concept aggregation for KT model, which enables KT model to retrive meta-knowledge from source domains, aggregate concepts from different domains and adapt to target domain.
In the knowledge tracing task, each question is typically associated with a few specific knowledge concepts, which is directly utilized by conventional knowledge tracing models. However, this approach is not applicable for domain generalization knowledge tracing tasks, as knowledge concepts can vary across different domains, and they may even share the same concept ID. Additionally, the features of interaction sequences from different knowledge tracing domains show significant variations. For example, in certain knowledge tracing tasks, students may practice a question repeatedly, while in other tasks, students will move on to the next question regardless of their performance on the current question. Consequently, the question sequences from different domains exhibit different levels of granularity, which poses challenges for knowledge tracing models to effectively analyze a student’s learning patterns.
To address the issue of granular differences, we propose concept aggregation for domain-specific knowledge concepts. This approach involves learning the domain-specific concept embedding and conducting $k$-means algorithm for embedding across all source domains. Subsequently, we replace concept embeddings with calculated centroid embedding and further train the model using concept embeddings.
Concept feature learning. We first train the KT model on all source domains. Each source domain learns their own concept embedding matrix $\textbf{E}_{i}=\{e_{j}\}_{j=1}^{n_{i}}$ (referred to Eq. (1)), while the remaining parameters of knowledge state encoder and decoder, namely $\theta_{enc}$ and $\theta_{dec}$ in KT model respectively, are shared across all domains. During this stage, we randomly sample data from each source domain. The model is trained by minimizing classification loss on all source domains, which can be formulated as:
@@ III-D Concept Aggregation
where $L_{\text{CE}}$ is the cross-entropy loss, $I^{i}_{j}$ is the $j$-th interaction sequence in the $i$-th source domain, and $R^{i}_{j}$ is the corresponding response to the next question.
Concept clustering. After first training on source domains, we obtain the concept embeddings $\{\textbf{E}_{i}\}_{i=1}^{N}$ from various domains. In this stage, we apply the $k$-means algorithm to all concept embeddings, dividing numerous concepts into $k$ clusters. This process provides KT model with a coarse-grained perspective of interaction sequences. Specifically, through the $k$-means procedure, we obtain a cluster assignment matrix A and the embeddings of each centroid $\textbf{E}_{c}$.
@@
where $\textbf{E}_{\text{cat}}$ concatenates all concept embeddings from $N$ source domains, and $n_{\text{e}}$ represents the total of the concepts. $\textbf{A}\in\{0,1\}^{k\times n_{\text{e}}}$ is the cluster assignment matrix where $\textbf{A}_{ij}=1$ means the $j$-th concept of $\textbf{E}_{\text{cat}}$ is assigned to the $i$-th centroid. $|C_{i}|$ denotes the number of concepts of the $i$-th cluster. $e_{c_{i}}$ denotes the embedding of $i$-th centroid and $\textbf{E}_{c}\in\mathbb{R}^{d\times k}$ represents the matrix of centroid embeddings.
Concept centroid refinement. Based on the results of concept clustering, we replaced the concept embeddings of each source domain with the embeddings of the cluster centroids to narrow the difference between the source domains. Accordingly, for domain x, the question embedding is rewritten as:
$\tilde{Q}$ $=A_{x}\cdot Q_{x},$ (20)
$e_{q_{t}}$ $=\frac{\sum_{i}\mathbb{I}(\tilde{Q}_{iq_{t}}=1)e_{i}}{\sum_{i}\mathbb{I}(\tilde{Q}_{iq_{t}}=1)},$ (21)
where $A_{x}\in\{0,1\}^{k\times n_{c_{x}}}$ is the cluster assignment matrix of domain $x$, where $\textbf{A}_{x_{ij}}=1$ means the $j$-th concept of $\textbf{E}_{x}$ is assigned to the $i$-th centroid. $\tilde{Q}$ represent the relationship between questions and centorids, where $\tilde{Q}_{ij}=1$ when j-th question is related to i-th centroid. $Q_{x}$ is the concept matrix of domain $x$. $e_{i}$ is the i-th vector in $\textbf{E}_{c}$, $\textbf{E}_{c}$ and A are calculated in concept clustering phase. Similar to Eq. (15), we randomly sample data from source domains and update the parameters ($\textbf{E}_{c}$, $\theta_{enc}$ and $\theta_{dec}$) in KT model by minimizing the classification loss.
The procedure of concept aggregation can be succinctly summarized as Algorithm 1.
Algorithm 1 Concept aggregation for source domain training
Input: Interactions from $N$ source domains $\{D_{s}^{i}\}_{i=1}^{N}$.
Output: Model parameters $\theta_{enc},\theta_{dec}$; Embedding $\textbf{E}_{c}$.
@@
12:  ///// / Concept clustering and concept centroid refinement
@@
III-E Generalization to Target Domain
For domain generalization of KT tasks, we delve into transferring the learned KT model to a designated target domain. Firstly, we initialize the concept embedding of the target domain and subsequently design a novel concept representation for the target domain which not only exhibits adaptability to the target domain but also maintains the centroid embeddings learned from the source domain to prevent overfitting.
Initialization for target embedding. We begin by initializing the concept embedding for the target domain, namely target embedding. Given that several concepts in the target domain remains unseen, the optimal approach for initializing the target embedding is to utilize the learned centroid embedding. For the target embedding, the formulation is as follows:
$e_{t_{i}}\in\{e_{c_{1}},e_{c_{2}},...,e_{c_{k}}\},$ (22)
$\textbf{E}_{t}=[e_{t_{1}},e_{t_{2}},...,e_{t_{n_{c}}}],$ (23)
where $e_{t_{i}}$ represents the $i$-th concept embedding in target domain, $1\leq i\leq n_{c_{t}}$. $n_{c_{t}}$ is the number of concepts in target domain. Initially, $e_{t_{i}}$ is randomly selected from the set of centroid embeddings. $\textbf{E}_{t}$ represents target embedding matrix.
Concept representation of target domain. Regarding the representation of concepts in the target domain, we aim for it to leverage the parameters learned during the concept aggregation phase effectively. Additionally, we seek its efficient adaptation to the target domain, enabling the model to swiftly transition with limited data while avoiding overfitting. For problems in target domain, we have:
$e_{q_{t}}=\frac{\sum_{i}\mathbb{I}(Q_{iq_{t}}=1)e_{i}}{\sum_{i}\mathbb{I}(Q_{iq_{t}}=1)},$ (24)
@@ III-E Generalization to Target Domain
where $e_{i}$ is the concept embedding retrived from $\textbf{E}_{t}$. $e_{c_{j}}$ denotes the $j$-th centroid embedding. $\lambda$ is a hyperparameter used to adjust the degree of freedom of the target concept embedding. $e_{target}$ is the final representation of $q_{t}$ in target domain that will be fed into KT model.
Target embedding adaptation. We fine-tune our KT model on target domain using the aforementioned concept representation. During this phase, only the target embedding $\textbf{E}_{t}$ is optimized, while other parameters including $\theta_{enc}$ and $\theta_{dec}$ are frozen, which is similar to Eq. (15).
In this way, our model will effectively adapt to target domain while fully utilizing its learned centroid embedding in concept aggregation. For clarification, the whole process of generalization is summarized in Algorithm 2.
Algorithm 2 Generalization to target domain
Input: Interactions from target domain $D_{t}$; Embedding $\textbf{E}_{c}$; Model parameters $\theta_{enc},\theta_{dec}$.
Output: Target embedding $\textbf{E}_{t}$.
1:  $\textbf{E}_{t}\leftarrow$ solution of Eqs. (22)-(23).  ///// / Initialization
2:  for $\text{iter}\leftarrow 1~{}\text{to}~{}\text{epoch}$ do
3:     $\mathcal{L}_{kt}\leftarrow 0$.
4:     for $\{I^{i},y^{i}\}\in D_{t}$ do
5:        $\{\textbf{M}_{q},\textbf{M}_{qr}\}\leftarrow$ solution of Eqs. (24)-(26).
6:        $\hat{y}\leftarrow\text{Dec}(\text{Enc}(\textbf{M}_{q},\textbf{M}_{qr}))$. ///// / Predict $\hat{y}$
7:        $\mathcal{L}_{kt}\leftarrow\mathcal{L}_{kt}+\mathcal{L}_{\text{CE}}(\hat{y},y^{i})$. ///// / Calculate the loss
8:     end for
9:     Update $\textbf{E}_{t}$ by minimizing $\mathcal{L}_{kt}$.
10:  end for

### v2

@@ 3.2 DGKT Architecture
where $\boldsymbol{0}=(0,0,...,0)\in\mathbb{R}^{d}$ is an all zero vector with the same dimension $d$ as $e_{q_{t}}$, and $\oplus$ is the concatenation operator.
@@
Knowledge state encoder using existing models. To encode the question embeddings and question-response embeddings into knowledge state, many off-the-shelf knowledge tracing models can be used as the knowledge state encoder:
$\{h_{t}\}=\text{Enc}(\textbf{M}_{q},\textbf{M}_{qr}),$ (3)
where the embedding matrices $\textbf{M}_{q}$ and $\textbf{M}_{qr}$ are encoded into obtain knowledge state $\{h_{t}\}$ at different timestamps.
Here we take DKT [28], SAINT [10] and AKT [15] as examples to illustrate how to transform embedding into knowledge state. Note that the architecture of SimpleKT [25] can also be used, but we will not elaborate on it due to its simplicity and similarity to SAINT and AKT.
For DKT [28], we use LSTM as knowledge state encoder to get student’s knowledge state:
$h_{t}=\text{LSTM}(e_{qr_{t}},h_{t-1}),$ (4)
where $e_{qr_{t}}$ denotes the $t$-th question-response embedding and, $h_{t}$ represents the student’s knowledge state at timestep $t$. The LSTM network takes knowledge state $h$ as its hidden state.
As for SAINT [10], it usually utilizes a Transformer architecture. The Transformer’s encoder is responsible for receiving the student’s question embedding matrix, while its decoder receives the student’s question-response embedding matrix along with the output from the encoder, which can be written by:
$\begin{split}\textbf{O}_{\{1:t\}}&=\text{Encoder}(\textbf{M}_{q\{1:t\}}),\\ h_{t}&=\text{Decoder}(\textbf{M}_{qr\{1:t-1\}},\textbf{O}_{\{1:t\}}),\end{split}$ (5)
where Encoder and Decoder represent the Transformer’s encoder and decoder, respectively.
Likewise, in the AKT [15] model, three attention-based modules are utilized, which are the question encoder, the knowledge encoder, and the knowledge retriever. The student’s knowledge state can be represented as:
$\begin{split}\textbf{X}_{\{1:t\}}&=\text{Encoder}_{q}(\textbf{M}_{q\{1:t\}}),\\ \textbf{Y}_{\{1:t-1\}}&=\text{Encoder}_{k}(\textbf{M}_{qr\{1:t-1\}}),\\ h_{t}&=\text{Decoder}(\textbf{X}_{\{1:t\}},\textbf{Y}_{\{1:t-1\}}),\end{split}$ (6)
where $\text{Encoder}_{q}$ represents the question encoder which produce contextualized representations of each question. $\text{Encoder}_{k}$ represents the knowledge encoder which produces contextualized representations of each question and response. Decoder represents the knowledge retriever which retrieves knowledge state.
In sum, the knowledge state encoder generates the knowledge state $h_{t}$, which will be sent through the knowledge state decoder.
Knowledge state decoder. We use a unified knowledge decoder to predict the probability of a student answering the next question correctly:
@@
where Dec represents the knowledge state decoder, and $W_{1}$, $W_{2}$ and $b_{1}$, $b_{2}$ denote the weights and the biases, respectively. Also, $\sigma(\cdot)$ is the sigmoid function.
Sequence Instance Normalization To reduce the distribution discrepancy from different domains, utilizing normalization method is a common solution. However, existing normalization methods (e.g., Batch Normalization [18] and Instance Normalization [33]) may not be applicable for sequential features in knowledge tracing tasks. For example, Instance Normalization normalizes all the information within a single channel, which can lead to the information leakage while predicting student’s performance at timestep $t$. This is because the student’s interactions after timestep $t$ should remain unseen in KT task. To tackle the dilemma, we design the Sequence Instance Normalization (SeqIN) method in our KT model to normalize the feature embeddings of sequential student interactions among domains. Since the normalization process takes into account the fact that the later interactions cannot be seen in the previous interactions, their feature embeddings at the current moment are normalized by only using the statistics of all the previous moments.
Given a sequential feature embedding matrix $\textbf{M}=(m_{1},m_{2},...,m_{n})$ where $m_{t}\in\mathbb{R}^{d}$ is the feature embedding at timestep $t$, we firstly calculate the mean $\mu_{t}$ and standard deviation $\sigma_{t}$ before timestep $t$:
$\mu_{t}=\frac{1}{t+1}(p+\sum_{i=1}^{t}m_{i}),$ (8)
$\sigma_{t}=\sqrt{\frac{(p-\mu_{t})^{2}+\sum_{i=1}^{t}(m_{i}-\mu_{t})^{2}}{t+1}},$ (9)
where $\mu_{t}$ and $\sigma_{t}$ denote the mean and standard deviation of the sequence $\{m_{1},m_{2},...,m_{t}\}$, and $p$ is a learnable padding vector. In consideration of the meaninglessness to compute the mean and standard deviation of $m_{1}$ itself, we add a padding vector $p$ in front of the original sequence, i.e., $\{p,m_{1},m_{2},...,m_{n}\}$, where $p$ is learned along with the model.
We then normalize the sequential feature embedding matrix M into $\tilde{\textbf{M}}$ using the calculated mean and standard deviation:
$\tilde{m}_{t}=\gamma(\frac{m_{t}-\mu_{t}}{\sigma_{t}})+\beta,$ (10)
$\tilde{\textbf{M}}=[\tilde{m}_{1},\tilde{m}_{2},...,\tilde{m}_{n}],$ (11)
where $\gamma,\beta\in\mathbb{R}^{d}$ are the learnable parameters. As seen, SeqIN normalizes $m_{t}$ by considering all previous feature embeddings up to time $t$ while remaining later feature embeddings unseen.
We apply SeqIN to aforementioned knowledge state encoders, including DKT, SAINT and AKT. For DKT, the SeqIN is directly used on students’ knowledge state $h$:
@@
where $\tilde{h}$ is the normalized students’ knowledge state.
As for SAINT, we apply SeqIN to intermediate features $o$:
$\tilde{o}=SeqIN(o),$ (13)
where $o$ is the output of Transformer’s encoder.
For AKT, we apply SeqIN to intermediate features $x$ and $y$ as follows:
@@
It is worthy noting that SeqIN can effectively aggregate the feature embeddings of student interactions from different source domains to reduce the domain gaps, which can be observed in the visualization experiment in Figure 8.
3.3 Concept Aggregation
To overcome significant distribution discrepancy among the source domains, we design the concept aggregation procedure for the KT model, which enables the retrieval of meta-knowledge from source domains through aggregating the concepts from diverse source domains, subsequently allowing for adaptation to the target domain.
In the knowledge tracing task, each question is typically associated with a few specific knowledge concepts, which is directly utilized by conventional knowledge tracing models. However, this approach is not suitable for multi-source domain adaptation knowledge tracing tasks, as knowledge concepts can differ significantly across various domains. On one hand, the feature embeddings of concepts from different domains exhibit significant distribution variations, as they belong to different education systems. On the other hand, these concepts may exhibit shared attributes, including difficulty, forgettability, among others.
To mitigate the challenge of substantial variations among concepts while maintaining their shared characteristics, we propose a concept aggregation strategy to generate concept prototypes. These prototypes represent the concepts from a unified perspective by aggregating similar domain-specific knowledge concepts. This concept aggregation process involves: 1) Concept feature learning for domain-specific concept embeddings, 2) Concept clustering by conducting $k$-means algorithm for the concept embeddings across all source domains, and subsequently, 3) Application of concept prototypes for re-representing question embeddings and model updating.
Concept feature learning. Initially, we train the KT model across all source domains. Each source domain generates its own concept embedding matrix, denoted as $\textbf{E}_{i}=[e_{1},e_{2},...,e_{n_{i}}]$, where $e_{i}\in\mathbb{R}^{d}$ represents the embedding of concept $i$. The remaining parameters of knowledge state encoder and decoder, $\theta_{enc}$ and $\theta_{dec}$, are shared across all domains within the KT model. During this phase, data from each source domain is randomly sampled to ensure that the model does not become biased towards the source domain with larger amounts of data. The model is trained by minimizing the classification loss across all source domains, which can be mathematically expressed as:
@@ 3.3 Concept Aggregation
where $L_{\text{CE}}$ is the cross-entropy loss, $I^{i}_{j}$ is the $j$-th interaction sequence in the $i$-th source domain, $\hat{Y^{i}_{j}}$ is the probability generated by the model, and $R^{i}_{j}$ is the ground truth, representing the corresponding response.
Concept clustering. After the initial training on source domains, we acquire the concept embeddings $\{\textbf{E}_{i}\}_{i=1}^{N}$ from various domains. In this phase, we apply the $k$-means algorithm to all the concept embeddings, grouping numerous concepts into $k$ concept prototypes. This procedure aggregates concepts that share common characteristics into a single concept prototype, even if these concepts originate from different domains. This approach helps reduce the distribution discrepancy between domains.
Specifically, through the $k$-means procedure, we first obtain a cluster assignment matrix A:
@@
where $\textbf{E}_{\text{cat}}$ is the concatenation of all concept embeddings from $N$ source domains. The matrix $\textbf{A}\in\{0,1\}^{k\times n_{\text{e}}}$ represents the cluster assignment matrix, where $\textbf{A}_{ij}=1$ indicates that the $j$-th concept from $\textbf{E}_{\text{cat}}$ is assigned to the $i$-th concept prototype. Then, we calculate the embedding of a concept prototype by averaging all concept embeddings in a cluster, and concate all concept prototypes to form a matrix, which can be written by:
@@
where $e_{c_{i}}$ denotes the embedding of $i$-th concept prototype and $\textbf{E}_{c}\in\mathbb{R}^{d\times k}$ represents the embedding matrix of concept prototypes. Here, $n_{\text{e}}$ represents the total of the concepts, and $|C_{i}|$ denotes the number of concepts of the $i$-th cluster.
Concept prototypes refinement. To reduce the differences between the source domains, we employ the concept prototypes to replace the original domain-specific concepts and to re-represent the question embeddings. Subsequently, these concept prototypes are further refined to update the model.
Specifically, for the $s$-th domain, we leverage the clustering assignment matrix $\textbf{A}_{s}$ of concepts and the concept matrix $\textbf{Q}_{s}$ to calculate the concept prototype matrix $\tilde{\textbf{Q}}$, i.e.,
$\tilde{\textbf{Q}}=\textbf{A}_{s}\cdot\textbf{Q}_{s},$ (21)
where $\tilde{\textbf{Q}}_{ij}=1$ indicates that the $j$-th question in the $s$-th domain is associated with the $i$-th concept prototype.
Then, the question embedding can be calculated by averaging all the embeddings of associated concept prototypes:
$\tilde{e_{q_{t}}}$ $=\frac{\sum_{i}\mathbb{I}(\tilde{\textbf{Q}}_{iq_{t}}=1)e_{c_{i}}}{\sum_{i}\mathbb{I}(\tilde{\textbf{Q}}_{iq_{t}}=1)},$ (22)
where $\tilde{e_{q_{t}}}$ is reformulated by using concept prototypes, as contrasted with Eq. (1). Subsequently, we update the embedding matrix ($\textbf{E}_{c}$) of the concept prototypes along with the model parameters ($\theta_{enc}$ and $\theta_{dec}$) within the KT model by minimizing the cross-entropy loss (as referenced in Eq. (16)).
Algorithm 1 Concept aggregation
Input: Interactions from $N$ source domains $\{D_{s}^{i}\}_{i=1}^{N}$ and target domain $D_{t}$.
Output: Model parameters $\theta_{enc},\theta_{dec},E_{c}$ and $E_{t}$.
@@
12:  ///// / Concept clustering and concept prototype refinement
@@
24:  ///// / Adaptation to target domain
25:  $\textbf{E}_{t}\leftarrow$ solution of Eq. (23).  ///// / Initialization
26:  for $\text{iter}\leftarrow 1~{}\text{to}~{}n_{epoch}$ do
27:     $\mathcal{L}_{kt}\leftarrow 0$.
28:     for $\{I^{i},y^{i}\}\in D_{t}$ do
29:        $\{\textbf{M}_{q},\textbf{M}_{qr}\}\leftarrow$ solution of Eqs. (25) and (26).
30:        $\hat{y}\leftarrow\text{Dec}(\text{Enc}(\textbf{M}_{q},\textbf{M}_{qr}))$. ///// / Predict $\hat{y}$
31:        $\mathcal{L}_{kt}\leftarrow\mathcal{L}_{kt}+\mathcal{L}_{\text{CE}}(\hat{y},y^{i})$. ///// / Calculate the loss
32:     end for
33:     Update $\textbf{E}_{t}$ by minimizing $\mathcal{L}_{kt}$.
34:  end for
After learning the domain-generalizable KT model, we delve into transferring the model to a new target domain, which remains unseen during the training phase. Firstly, we randomly initialize the concept embeddings for the target domain. Following this, we design a novel question representation for the target domain that not only preserve the specific concepts of the target domain but also demonstrates adaptability to the concept prototypes learned from the source domains, thereby preventing overfitting.
Initialization for target concept embeddings. We begin by initializing the concept embeddings for the target domain, namely target concept embeddings. Given that concepts in the target domain remains unseen, a feasible solution for initializing the target concept embeddings is to randomly select ones from the concept prototypes to form the matrix $\textbf{E}_{t}$ of target concept embeddings, which can be formulated as:
$\textbf{E}_{t}=[e_{t1},e_{t2},...,e_{t{n_{c}}}],$ (23)
where $e_{ti}$ represents the $i$-th target concept embedding ($1\leq i\leq n_{c}$), and $n_{c}$ is the number of concepts in target domain.
Question representation of target domain. We aim to effectively leverage the concept prototypes learned from various source domains to represent the question embeddings of the target domain. Firstly, as in Eq. (1), we calculate the concept-based question embedding $e_{q_{t}}$ as follows:
$e_{q_{t}}=\frac{\sum_{i}\mathbb{I}(\textbf{Q}_{iq_{t}}=1)e_{ti}}{\sum_{i}\mathbb{I}(\textbf{Q}_{iq_{t}}=1)},$ (24)
where $e_{ti}$ is the concept embedding retrieved from $\textbf{E}_{t}$. Then, given the potential risk of overfitting due to limited student interactions in the target domain, we further constrain the question embedding to move closer to its nearest concept prototype. Thus, in the target domain, the question embedding $e_{target}$ can finally be calculated as:
@@
where $e_{c_{j}}$ represents the nearest concept prototype to $e_{q_{t}}$, $\lambda$ is a hyperparameter used to control the degree of influence that the target concept embedding has over the final question embedding. This ensures that the question embedding $e_{target}$ not only captures the relevant features of the target domain but also remains aligned with the generalized knowledge encoded in the concept prototypes, reducing the likelihood of overfitting.
Adaptation of target concept embeddings. Our KT model is finetuned on the target domain by utilizing the question embedding $e_{target}$ within the cross-entropy loss. During this phase, optimization is confined to the target concept embedding $\textbf{E}_{t}$, with all the other parameters including $\theta_{enc}$ and $\theta_{dec}$, respectively, remaining fixed. In this way, our model will effectively adapt to the target domain while leveraging the concept prototypes embeddings learned during concept aggregation. For further clarity, the whole process of concept aggregation detailed in Algorithm 1.
In concept aggregation, the space complexity of our feature embedding module is expressed as $O\left(\sum_{i}^{n}n_{c_{i}}+n_{c_{t}}+k\right)$, where $n_{c_{i}}$ denotes the number of concepts in the source domain $i$, $n_{c_{t}}$ denotes the number of concepts in the target domain, and $k$ is the number of concept prototypes involved in concept aggregation.
3.4 Relation-Aware Knowledge Encoder
Although the proposed DGKT framework enables existing knowledge tracing (KT) models to achieve domain generalizability and enhances their performance in new domains, these models are not inherently designed for cross-domain tasks. A significant challenge impeding the effective transfer of these models to new domains is the limited availability of training data, which hampers the model’s ability to fully exploit the information embedded in students’ interaction sequences.
To mitigate this limitation, we propose a domain-generalizable relation-aware knowledge encoder (RA-Encoder), explicitly designed to address the challenge of multi-source domain adaptation in knowledge tracing. This encoder leverages an attention mechanism to capture the relationships between questions and concepts across different timesteps. In this section, we provide a detailed exposition of the RA-Encoder, including its architecture and the DGrKT model utilizing RA-Encoder.
Relation-Aware Attention Encoder In the Relation-Aware Attention Encoder (RA-Encoder), the attention value $\alpha$ is modulated according to the relation matrix between any two distinct timesteps, as illustrated in Fig. 5. Our underlying assumption is that the more similar a question a student previously answered is to the current question, the more valuable the information from that previous time becomes. Therefore, we have categorized the relationships between different timesteps into three distinct types to facilitate the learning process of the relation-aware attention encoder.